In pilot plant for iodine extraction, 4,000 L of chlorine gas at 1.5 atm and 30oC reacts with 1,500 L of seawater with 1.3 X 10-3 M iodide ion

1) How many moles of chlorine gas are available to react with the seawater? Show your work to receive full credit.

2) How many moles of iodide ions (I-) are available to react with the chlorine gas? Show your work to receive full credit.

3) How many grams (g) of solid iodine (I2) can the pilot plant theoretically produce? Show your work to receive full credit.
1
Expert's answer
2020-12-03T13:54:13-0500

1) pV=nRT. n=pV/RT = 1.5 x 4/0.082 x 303 = 0.24 mol of chlorine.

2) Cm=n/V. n= Cm x V = 1.3 x 10^-3 x 1.5 = 0,00195 mol of iodide ions.

3) Cl2 + 2 HJ= 2HCl + J2.

0.00195 x 254/2 = 0,24765 g of solid iodine.